OpenStudy (anonymous):

Bill's upstream canoe trip took 2 hours longer than the return trip. If he averaged 2 mph on the upstream trip and 6 mph on the return trip, then how long was the upstream trip?

OpenStudy (sandra):

1. u = d + 2 And given they're the same distance: 2. u*2 = d*6 now substitute d+2 for u in the second equation: (d+2)*2 = d*6 2d + 4 = 6d 4 = 6d-2d 4 = 4d d= 1. The downstream trip took 1 hour

OpenStudy (anonymous):

thank you